ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 116, 2000

Agaoka, Y. Left invariant Poisson structures on classical noncompact simple Lie groups

189

Appell, J., C. Franchetti and E. M. Semenov Estimates for the Jung constant in Banach lattices

171

Arnold, L., D. C. Nguyen and V. I. Oseledets The essential range of a nonabelian cocycle is not a cohomology invariant

71

Barreira, L. and J. Schmeling Sets of “non-typical” points have full topological entropy and full Hausdorff dimension

29

Bell, H. E. See A. A. Klein

Esterle, J. and R. Gay Product of hyperfunctions on the circle  271

Franchetti, C. See J. Appell Gay, R. See J. Esterle Higgins, P. M. and S. W. Margolis Finite aperiodic semigroups

with commuting idempotents and generalizations

367

Kaliman, S. Uniform Zariski’s theorem on fundamental groups  323
Kaufman, R. On smooth norms and analytic sets  21

Klein, A. A. and H. E. Bell On the multiplicative semigroup of a ring

249

Leung, K. H. Forms and Baer ordered ∗-fields  1
Ling, S. Component group of the p-new subvariety of J_{0}(Mp)  117

López-Bautista, P. R. and G. D. Villa-Salvador On the Galois module structure of semisimple holomorphic differentials

345

Margolis, S. W. See P. M. Higgins Masumoto, M. Hyperbolic lengths and conformal embeddings of

Riemann surfaces

77

Nguyen, D. C. See L. Arnold Oseledets, V. I. See L. Arnold Saussol, B. Absolutely continuous invariant measures for multi-

dimensional expanding maps

223

Schmeling, J. See L. Barreira

Schoutens, H. Bounds in cohomology  125

Semenov, E. M. See J. Appell

Shelah, S. The generalized continuum hypothesis revisited  285

Steinberger, T. Hausdorff dimension of attractors for two dimensional Lorenz transformations

253

Villa-Salvador, G. D. See P. R. López-Bautista

Watanabe, T. A comparison of automorphic L-functions in a theta series lifting for unitary groups 93